# Glauber predictions for oxygen and neon collisions at the LHC

### v2

@@ IV Nuclear density profiles for oxygen and neon
Instead of using average $\rho$ distributions, explicit calculations of the nucleon configuration is desired to account for inter-nucleon correlations, in particular important for small systems. Previous work Lim:2018huo used an ab-initio quantum Monte Carlo calculation of the Oxygen nucleon configurations with two- and three-body local chiral potentials at next-to-next-to-leading order Lonardoni:2017hgs , which were available in TGlauberMC (since v2.7, named “O”, see Tab. 1). Recently, Oxygen and Neon configurations where derived within the framework of Nuclear Lattice Effective Field Theory (NLEFT) and the ab-initio Projected Generator Coordinate Method (PGCM), and used as input for the initial conditions in Trajectum Giacalone:2024luz . The NLEFT framework is well suited to probe collective phenomena in the ground states of nuclei Lu:2018bat . Similarly to Summerfield:2021oex , the positions of the nucleons were traced using the pinhole Elhatisari:2017eno algorithm. In this way, the NLEFT configurations, which can be obtained keeping track of positive and negative weights, and with resolving ambiguities introduced by the underlying lattice periodicity, capture short-range correlations without explicitly enforcing $\alpha$-clusters. In the PGCM approach Frosini:2021ddm ; Huther:2019ont , a set of mean-field states with different quadrupole and octupole deformations is first generated to explore a range of possible nuclear shapes. Each of these intrinsic configurations is then projected onto states with physical particle number and angular momentum to restore the symmetries broken at the mean-field level. A collective wavefunction is constructed by mixing the projected states through a variational procedure that minimizes the total energy of the system. The final nucleon configurations are then constructed with or without explicitly enforcing $\alpha$-clustering. A summary of the physically meaningful nucleon configurations obtained by the NLEFT and PGCM methods is given in Tab. 2. For more details, see the extra material provided in Ref. Giacalone:2024luz .
@@ V Cross section results
To compute the expected OO, NeNe and pO geometrical cross sections at 5.36 and 9.62 TeV, needs the values for $\sigma_{\rm NN}$ at the respective nucleon–nucleon center-of-mass beam energy as input to the MCG calculation. The $\sigma_{\rm NN}$ can be obtained from the parametrization, discussed in dEnterria:2020dwq ,
@@
The resulting cross sections for OO and NeNe collisions at 5.36 TeV are shown in Fig. 4, for $d_{\rm min}=0.4$ fm and $\omega=0$, i.e. in the hard-sphere approximation, used in previous works Loizides:2017ack ; dEnterria:2020dwq . The cross sections averaged over all profiles discussed above are $\sigma_{\rm HS}=1313\pm 44$ mb and $\sigma_{\rm HS}=1678\pm 30$ mb, for OO and NeNe collisions, respectively. The above uncertainties are obtained by using the nuclear profiles as described in Sec. IV. Instead, in case of oxygen, varying the two parameters of the “Oho2” model, within the reported uncertainties would results in a relative uncertainty below 1%. The resulting cross section for pO collisions at 9.62 TeV are shown in extra Fig. 16, leading to $\sigma_{\rm HS}=466\pm 9$ mb when averaging over the various oxygen density profiles.
Recentering (discussed above) does not affect the calculated cross section, neither does changing $d_{\rm min}$ to a value lower than 0.4 fm. Increasing $d_{\rm min}$ beyond 0.4 fm, however starts to affect the calculated cross section (e.g. for $d_{\rm min}=0.8$ fm the increase is nearly 2.5%). However using $d_{\rm min}=0.8$ fm significantly distorts the charged density, which is expected giving that the average minimal separation is less than 0.8 fm (evaluated using the nucleon configurations Lonardoni:2017hgs named “O” in TGlauberMC). Variation of $\sigma_{\rm NN}$ by $\pm 1.2$ mb, leads to a change of less than 1%.
@@
The choice of $\omega=0$ is not really justified since particle production in pp collisions strongly depends on the nucleon–nucleon impact parameter distribution, exemplified in the development of the multi-parton picture of the PYTHIA Sjostrand:2017cdm generator to describe data obtained at hadron colliders in the TeV scale. However, as discussed in Sec. III there is a large parameter space beyond the hard-sphere approximation. In particular, the calculated nucleus–nucleus cross section depends quite significantly on the nucleon–nucleon collision probability, as already discussed in Ref. Nijs:2022rme . The dependence of the OO cross section on $\omega$ is shown in Fig. 5. The cross section is found to increase rather linearly with $\omega$ with a slope of about $161$ mb, i.e. the relative increase to the hard-sphere approximation is about $12.3\%\cdot\omega$. In addition, results from HIJING, PYTHIA and TRENTO with $w=0.5$ are indicated, which roughly span the values for $\omega$ between $0.3$ and $0.6$.
@@
Table 3: Measured and calculated cross sections for PbPb collisions at 2.76 TeV, and pPb and PbPb collisions at 5.02 TeV. The calculations using the hard-sphere approximation are from Loizides:2017ack . The new results using the $\Gamma$-parametrization for the nucleon–nucleon overlap function with $\omega=0.3$ are obtained in this work as motivated in the text.
System Measurement Glauber (HS) Loizides:2017ack Glauber ($\omega=0.3$)
PbPb, $\sqrt{s_{\rm NN}}=2.76$ TeV $\sigma_{\rm PbPb}=7.7\pm 0.6$ b ALICE:2012aa $\sigma_{\rm PbPb}^{\rm HS}=7.55\pm 0.15$ b $\sigma_{\rm PbPb}^{\rm MC}=7.70\pm 0.30$ b
PbPb, $\sqrt{s_{\rm NN}}=5.02$ TeV $\sigma_{\rm PbPb}=7.67\pm 0.25$ b ALICE:2022xir $\sigma_{\rm PbPb}=7.62\pm 0.15$ b $\sigma_{\rm PbPb}^{\rm MC}=7.80\pm 0.29$ b
pPb, $\sqrt{s_{\rm NN}}=5.02$ TeV $\sigma_{\rm pPb}=2.06\pm 0.03\pm 0.1$ b CMS:2015nfb $\sigma_{\rm pPb}^{\rm HS}=2.09\pm 0.03$ b $\sigma_{\rm pPb}^{\rm MC}=2.13\pm 0.07$ b
$\sigma_{\rm pPb}=2.10\pm 0.07$ b ALICE:2014gvw
@@
To help deciding on a value for $\omega$, we list existing cross section data at the LHC in Tab. 3 together with earlier calculations using the hard-sphere approximation Loizides:2017ack . As can be seen, there is not a lot of room to increase the cross section beyond the hard-sphere approximation. Quite the contrary, the calculated cross sections are almost perfectly agreeing with the measurements. The recent Bayesian analysis Nijs:2022rme mentioned earlier, which includes a wealth of data points different from cross sections, obtained $0.4<w<0.5$ for the TRENTO model, which is similar to $\omega=0.3$ case (see Fig. 5 and Fig. 6). Indeed, the $\chi^{2}/n_{\rm dof}$ between all the available data points and the cross section dependence on $\omega$ extracted by the linear fits changes by about 1 for $\omega=0$ to $0.3$, as the associated increase is of similar size as the systematic uncertainty quoted for the hard-sphere based calculation. It hence is reasonable to increase the predicted cross sections by the associated relative amount (between 1.5% and 2%), and assign the same value as additional uncertainty. In this way, also the HIJING overlap profile, which qualitatively describes the centrality dependence of particle production in many collision systems ALICE:2014xsp ; Loizides:2017sqq ; ALICE:2018ekf ; dEnterria:2020dwq , is adequately covered. Using this prescription, the predicted cross sections for $\omega=0.3$ are given in Tab. 3, together with the values using the hard-sphere approximation.
VI Initial-conditions related distributions
@@
Figure 9: Ratio of participant eccentricity for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2.
In the following, we discuss the number of participants ($N_{\rm part}$) typically needed to compare measured quantities across systems and number of collisions ($N_{\rm coll}$) typically used to quantify the nuclear modification factor. Average $N_{\rm coll}$ can be obtained using $N_{\rm coll}=A\,B\,\sigma_{\rm NN}/\sigma_{\rm AB}^{\rm MC}$ leading to $12.8\pm 0.8$, $15.8\pm 1.1$, $2.49\pm 0.12$ for OO and NeNe at 5.36 TeV and pO at 9.62 TeV (see Tab. 4). In case of pO (or pPb) collisions, it follows that average $N_{\rm part}=\langle N_{\rm coll}\rangle+1$.
Figure 7 shows the $N_{\rm coll}$ and $N_{\rm part}$ distributions versus the impact parameter $b$ for OO and NeNe collisions at 5.36 TeV computed with TGlauberMC ($d_{\rm min}=0.4$fm, $\omega=0.3$) using the nuclear density profiles discussed in Sec. IV. In more central collisions (smaller impact parameter) the $N_{\rm coll}$ values differ with respect to the average by 5% for oxygen and up 10% for neon. In peripheral collisions (larger impact parameter) the differences increase to 10% and 15%, respectively, and a grouping into different types of underlying nuclear density profiles (Wood-Saxon, NLEFT, PGCM) becomes apparent. For $N_{\rm part}$ the differences are smaller, on the level of few percent in central collisions, reaching 5% for oxygen and 10% for neon in peripheral collisions. In the actual measurements, the observables will be measured in intervals of multiplicity and/or energy, which for large nuclei are well correlated with centrality. The intervals then will be mapped to $N_{\rm part}$ (and $N_{\rm coll}$ with different procedures summarized in Ref. Miller:2007ri ; dEnterria:2020dwq , which rely on MCG calculations. The process requires the data and calculations to be matched at the so-called anchor point. In PbPb collisions the anchor point can be determined rather precisely to the 1% level, and up to 90% peripheral events ALICE:2013hur . For OO (and NeNe) collisions, this may not be achievable, due to the much larger fluctuations and the unknowns of the underlying nuclear density profile.
@@
Figure 9 shows the ratio of participant eccentricity for NeNe collisions to that for OO collisions at 5.36 TeV for the ratio of the averages (including the standard deviation as uncertainty) as well as direct ratios from the nuclear densities listed in Tab. 2. As can be seen, the average ratio rises with increasing impact parameter, reaching about 1.05 in central collision. On the other hand, the NLEFT and PGCM profiles reach values significantly larger values up to 1.10–1.15. Assuming an hydrodynamical expansion of a nearly inviscid fluid, $\varepsilon_{2}\propto v_{2}$ holds to first order Noronha-Hostler:2015dbi , the ratio of $v_{2}$ between oxygen and neon is predicted to be of similar size. Indeed, a detailed calculation Giacalone:2024luz expects the ratio to be $1.15$ to within $\pm 3\%$ for most central collisions, since most theoretical uncertainties cancel out in the $v_{2}$ ratio.
@@
Figure 11: Ratio of $\epsilon_{3}$ (left panel) and $\epsilon_{4}$ (right panel) for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2.
For completeness, the distributions for $\varepsilon_{3}$ and $\varepsilon_{4}$ versus impact parameter for OO and NeNe collisions at 5.36 TeV are shown in Fig. 10 using the nuclear density profiles listed in Tab. 2. Both exhibit an increasing trend from central to peripheral collisions as known from larger nuclei collision systems. Variations of up to 10% relative to the average are observed as a result of the differences between the nuclear density profiles. Figure 11 shows the ratios of $\varepsilon_{3}$ and $\varepsilon_{4}$ for NeNe collisions to that for OO collisions at 5.36 TeV for the ratio of the averages (including the standard deviation as uncertainty) as well as direct ratios from the nuclear densities listed in Tab. 2. The ratio of $\varepsilon_{3}$ between NeNe and OO exhibits a modest decrease with decreasing centrality, turning into a strong rise for most peripheral collisions. The ratio of $\varepsilon_{4}$ between NeNe and OO decreases with decreasing centrality by about 15% with a modest rise towards most peripheral collisions.

### v3

@@ IV Nuclear density profiles for oxygen and neon
TR_NEON_V13 PGCM, constraints imposed on projected states, without explicit $\alpha$ clustering ($C=0$)
TR_NEON_V14 PGCM, constraints imposed on projected states, with explicit $\alpha$ clustering ($C=1$)
Instead of sampling nucleon positions from average $\rho$ distributions, explicit calculations of the nucleon configuration based on the many-body wave function, is desired to account for inter-nucleon correlations, in particular important for small and deformed nuclei Dobaczewski et al. (2025). In this case, an explicit $d_{\rm min}$ criterion between nucleons should not be imposed. Previous work Lim et al. (2019) used an ab-initio quantum Monte Carlo calculation of the Oxygen nucleon configurations with two- and three-body local chiral potentials at next-to-next-to-leading order Lonardoni et al. (2018), which were available in TGlauberMC (since v2.7, named “O”, see Tab. 1). Recently, Oxygen and Neon configurations where derived within the framework of Nuclear Lattice Effective Field Theory (NLEFT) and the ab-initio Projected Generator Coordinate Method (PGCM), and used as input for the initial conditions in Trajectum Giacalone and others (2025). The NLEFT framework is well suited to probe collective phenomena in the ground states of nuclei Lu et al. (2019). Similarly to Summerfield et al. (2021), the positions of the nucleons were traced using the pinhole Elhatisari et al. (2017) algorithm. In this way, the NLEFT configurations, which can be obtained keeping track of positive and negative weights, and with resolving ambiguities introduced by the underlying lattice periodicity, capture short-range correlations without explicitly enforcing $\alpha$-clusters. In the PGCM approach Frosini et al. (2022); Hüther et al. (2020), a set of mean-field states with different quadrupole and octupole deformations is first generated to explore a range of possible nuclear shapes. Each of these intrinsic configurations is then projected onto states with physical particle number and angular momentum to restore the symmetries broken at the mean-field level. A collective wavefunction is constructed by mixing the projected states through a variational procedure that minimizes the total energy of the system. The final nucleon configurations are then constructed with or without explicitly enforcing $\alpha$-clustering, denoted as “PGCM, C=1C=1” and “PGCM, C=0C=0”, respectively. A summary of the physically meaningful nucleon configurations obtained by the NLEFT and PGCM methods is given in Tab. 2. For more details, see the extra material provided in Ref. Giacalone and others (2025).
@@ V Cross section results
To compute the expected OO, NeNe and pO geometrical cross sections at 5.36 and 9.62 TeV, one needs the values for $\sigma_{\rm NN}$ at the respective nucleon–nucleon center-of-mass beam energy as input to the MCG calculation. The $\sigma_{\rm NN}$ can be obtained from the parametrization, discussed in d’Enterria and Loizides (2021),
@@
The resulting cross sections for OO and NeNe collisions at 5.36 TeV are shown in Fig. 4, for $d_{\rm min}=0.4$ fm and $\omega=0$, i.e. in the hard-sphere approximation, used in previous works Loizides et al. (2018); d’Enterria and Loizides (2021). The cross sections averaged over all profiles discussed above are $\sigma_{\rm HS}=1313\pm 44$ mb and $\sigma_{\rm HS}=1678\pm 30$ mb, for OO and NeNe collisions, respectively. The above uncertainties are obtained by using the nuclear profiles as described in Sec. IV. Instead, in case of oxygen, varying the two parameters of the “Oho2” model, within the reported uncertainties would results in a relative uncertainty below 1%. The resulting cross section for pO collisions at 9.62 TeV are shown in extra Fig. 17, leading to $\sigma_{\rm HS}=466\pm 9$ mb when averaging over the various oxygen density profiles.
Recentering (discussed above) does not affect the calculated cross section, neither does changing $d_{\rm min}$ to a value lower than 0.4 fm. Increasing $d_{\rm min}$ beyond 0.4 fm, however starts to affect the calculated cross section (e.g. for $d_{\rm min}=0.8$ fm the increase is nearly 2.5%). However using $d_{\rm min}=0.8$ fm significantly distorts the charged density, which is expected given that the average minimal separation is less than 0.8 fm (evaluated using the ab-initio nucleon configurations Lonardoni et al. (2018) named “O” in TGlauberMC). Variation of $\sigma_{\rm NN}$ by $\pm 1.2$ mb, leads to a change of less than 1%.
@@
The choice of $\omega=0$ is not really justified since particle production in pp collisions strongly depends on the nucleon–nucleon impact parameter distribution, exemplified in the development of the multi-parton picture of the PYTHIA Sjöstrand (2018) generator to describe data obtained at hadron colliders in the TeV scale. However, as discussed in Sec. III there is a large parameter space beyond the hard-sphere approximation. In particular, the calculated nucleus–nucleus cross section depends quite significantly on the nucleon–nucleon collision probability, as already discussed in Ref. Nijs and van der Schee (2022a). The dependence of the OO cross section on $\omega$ is shown in Fig. 5. The cross section is found to increase rather linearly with $\omega$ with a slope of about $161$ mb, i.e. the relative increase to the hard-sphere approximation is about $12.3\%\cdot\omega$. In addition, results from HIJING, PYTHIA and TRENTO with $w=0.5$ are indicated, which roughly span the values for $\omega$ between $0.3$ and $0.6$.
@@
Table 3: Measured and calculated cross sections for PbPb collisions at 2.76 TeV, and pPb and PbPb collisions at 5.02 TeV. The calculations using the hard-sphere approximation are from Loizides et al. (2018). The new results using the $\Gamma$-parametrization for the nucleon–nucleon overlap function with $\omega=0.3$ are obtained in this work as motivated in the text.
System Measurement Glauber (HS) Loizides et al. (2018) Glauber ($\omega=0.3$)
PbPb, $\sqrt{s_{\rm NN}}=2.76$ TeV $\sigma_{\rm PbPb}=7.7\pm 0.6$ b Abelev and others (2012) $\sigma_{\rm PbPb}^{\rm HS}=7.55\pm 0.15$ b $\sigma_{\rm PbPb}^{\rm MC}=7.70\pm 0.30$ b
PbPb, $\sqrt{s_{\rm NN}}=5.02$ TeV $\sigma_{\rm PbPb}=7.67\pm 0.25$ b Acharya and others (2024a) $\sigma_{\rm PbPb}=7.62\pm 0.15$ b $\sigma_{\rm PbPb}^{\rm MC}=7.80\pm 0.29$ b
pPb, $\sqrt{s_{\rm NN}}=5.02$ TeV $\sigma_{\rm pPb}=2.06\pm 0.03\pm 0.1$ b Khachatryan and others (2016) $\sigma_{\rm pPb}^{\rm HS}=2.09\pm 0.03$ b $\sigma_{\rm pPb}^{\rm MC}=2.13\pm 0.07$ b
$\sigma_{\rm pPb}=2.10\pm 0.07$ b Abelev and others (2014)
@@
To help deciding on a value for $\omega$, existing cross section data at the LHC are provided in Tab. 3 together with earlier calculations using the hard-sphere approximation Loizides et al. (2018). As can be seen, there is not a lot of room to increase the cross section beyond the hard-sphere approximation. Quite the contrary, the calculated cross sections are almost perfectly agreeing with the measurements. The recent Bayesian analysis Nijs and van der Schee (2022a) mentioned earlier, which includes a wealth of data points different from cross sections, obtained $0.4<w<0.5$ for the TRENTO model, which is similar to $\omega=0.3$ case (see Fig. 5 and Fig. 6). Indeed, the $\chi^{2}/n_{\rm dof}$ between the available data points and the $\omega$-dependent cross section obtained from the linear fits is about $0.6$ for $\omega=0.3$, and the corresponding increase is comparable in size to the systematic uncertainty quoted for the hard-sphere–based calculation. It hence is reasonable to increase the predicted cross sections by the associated relative amount (between 1.5% and 2%), and assign the same value as additional uncertainty. In this way, also the HIJING overlap profile, which qualitatively describes the centrality dependence of particle production in many collision systems Adam and others (2015); Loizides and Morsch (2017); Acharya and others (2019a); d’Enterria and Loizides (2021), is adequately covered. Using this prescription, the predicted cross sections for $\omega=0.3$ are given in Tab. 3, together with the values using the hard-sphere approximation.
VI Initial-condition related distributions
@@
Figure 9: Ratio of participant eccentricity versus centrality for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2.
Figure 10: Ratio of participant eccentricity versus centrality for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2. Gaussian smearing with $p=1$ (left) and $p=0$ (right panel) for the nuclear overlap is used as explained in the text.
In the following, the number of participants ($N_{\rm part}$) typically needed to compare measured quantities across systems and number of collisions ($N_{\rm coll}$) typically used to quantify the nuclear modification factor are discussed. Average $N_{\rm coll}$ can be obtained using $N_{\rm coll}=A\,B\,\sigma_{\rm NN}/\sigma_{\rm AB}^{\rm MC}$ leading to $12.8\pm 0.8$, $15.8\pm 1.1$, $2.49\pm 0.12$ for OO and NeNe at 5.36 TeV and pO at 9.62 TeV (see Tab. 4). In case of pO (or pPb) collisions, it follows that average $N_{\rm part}=\langle N_{\rm coll}\rangle+1$.
Figure 7 shows the $N_{\rm coll}$ and $N_{\rm part}$ distributions versus the impact parameter $b$ for OO and NeNe collisions at 5.36 TeV computed with TGlauberMC ($d_{\rm min}=0.4$fm, $\omega=0.3$) using the nuclear density profiles discussed in Sec. IV. In more central collisions (smaller impact parameter) the $N_{\rm coll}$ values differ with respect to the average by 5% for oxygen and up 10% for neon. In peripheral collisions (larger impact parameter) the differences increase to 10% and 15%, respectively, and a grouping into different types of underlying nuclear density profiles (Wood-Saxon, NLEFT, PGCM) becomes apparent. For $N_{\rm part}$ the differences are smaller, on the level of few percent in central collisions, reaching 5% for oxygen and 10% for neon in peripheral collisions. In the actual measurements, the observables will be measured in intervals of multiplicity and/or energy, which for large nuclei are well correlated with centrality. The intervals then will be mapped to $N_{\rm part}$ (and $N_{\rm coll}$) with different procedures summarized in Ref. Miller et al. (2007); d’Enterria and Loizides (2021), which rely on MCG calculations. The process requires the data and calculations to be matched at the so-called anchor point. In PbPb collisions the anchor point can be determined rather precisely to the 1% level, and up to 90% peripheral events Abelev and others (2013). For OO (and NeNe) collisions, this may not be achievable, due to the much larger fluctuations and the unknowns of the underlying nuclear density profile.
@@
Figure 9 shows the ratio of participant eccentricity for NeNe collisions to that for OO collisions at 5.36 TeV versus centrality (estimated from the respective impact parameter distributions) for the ratio of the averages (including the standard deviation as uncertainty) as well as direct ratios from the nuclear densities listed in Tab. 2. As can be seen, the average ratio rises with increasing impact parameter, reaching about 1.05 in central collision. On the other hand, the NLEFT and PGCM profiles reach values significantly larger values up to 1.10–1.15. Assuming an hydrodynamical expansion of a nearly inviscid fluid, $\varepsilon_{2}\propto v_{2}$ holds to first order Noronha-Hostler et al. (2016), the ratio of $v_{2}$ between oxygen and neon is predicted to be of similar size. Indeed, a detailed calculation Giacalone and others (2025) expects the ratio to be $1.15$ to within $\pm 3\%$ for most central collisions, since most theoretical uncertainties cancel out in the $v_{2}$ ratio.
@@
By construction, the participant eccentricity approaches unity in the most peripheral collisions when it is calculated directly from the participant positions, which in the limiting case may consist of only two participants. This behavior was first investigated in Ref. Alver and others (2008), which demonstrated for AuAu collisions that applying (Gaussian) smearing to the participant positions primarily affects the participant eccentricity in peripheral events (up to approximately $N_{\rm part}\sim 30$). For small systems, however, this range effectively covers the entire domain of interest, and smearing therefore has a substantial impact. To study whether smearing has an effect on the predicted ratio $\varepsilon_{2}^{\rm NeNe}/\varepsilon_{2}^{\rm OO}$, one commonly relies on the generalized mean for the nuclear overlap function defined as $\left(\frac{(T_{A}^{p}+T_{B}^{p})}{2}\right)^{\frac{1}{p}}$ Moreland et al. (2015) which for $p=1$ reduces to the arithmetic mean and for $p=0$ to the geometric mean (also known as TRENTO initial conditions). As shown in Fig. 10, in both cases the ratio approaches similar large values towards most central collisions as without Gaussian smearing. However, the trend with decreasing centrality differs, in particular for the TRENTO ($p=0$) condition, which increases with decreasing centrality.
For completeness, the distributions for $\varepsilon_{3}$ and $\varepsilon_{4}$ versus impact parameter for OO and NeNe collisions at 5.36 TeV are shown in Fig. 11 using the nuclear density profiles listed in Tab. 2. Both exhibit an increasing trend from central to peripheral collisions as known from larger nuclei collision systems. Variations of up to 10% relative to the average are observed as a result of the differences between the nuclear density profiles. Figure 12 shows the ratios of $\varepsilon_{3}$ and $\varepsilon_{4}$ for NeNe collisions to that for OO collisions at 5.36 TeV for the ratio of the averages (including the standard deviation as uncertainty) as well as direct ratios from the nuclear densities listed in Tab. 2. The ratio of $\varepsilon_{3}$ between NeNe and OO exhibits a modest decrease with decreasing centrality, turning into a strong rise for most peripheral collisions. The ratio of $\varepsilon_{4}$ between NeNe and OO changes only little with centrality, most notably with a decreasing trend with decreasing centrality for the NLEFT calculation by about 5 percent.
Figure 12: Ratio of $\epsilon_{3}$ (left panel) and $\epsilon_{4}$ (right panel) versus centrality for NeNe collisions to that for OO collisions at 5.36 TeV using TGlauberMC with $\omega=0.3$ for the ratio of the averages (standard deviation is shown as uncertainty per point), example ratio using “Ne2” and “Oho2” parameterizations as well as the ratios obtained using the nuclear densities listed in Tab. 2.